Marine biologists have noticed that the color of the outermost growth band on a clam tends to be related to the time of the year in which the clam dies. A biologist conducted a small investigation of whether this is true for the species Protothaca staminea. She collected a sample of 136 clam shells from this species and classified them according to the month that the clam died (February or March) and the color of the outermost growth band (clear, dark, or unreadable). The data are given in the Minitab output below.

-------------------------------------------------------------
Rows: Month Columns: Color
1 2 3 All
1 ? 28 ? 62
? ? ?
2 6 ? 3 ?
? ? ?
All ? ? 15 ?
Cell Contents: Count
Expected count
Chi-Sq = 2.377, DF = 2, P-Value = 0.305
-------------------------------------------------------------

If a chi-square analysis were performed on the above data, what would be the value in cell (1,1) of the expected table?
Answer #1

from above:

Contingency table should be:

Observed column 1 column 2 column 3 Total
row 1 28 62
row 2 6 3
total 0 35 15

cell (1,3) =15-3 =12

therefore cell (1,1) =62-28-12 =22

therefore value in cell (1,1) of the expected table =row total*column total/grand total

=28*62/136 =12.7647